How many liters of an 11% alcohol solution should be mixed with 60 liters of a 5% alcohol solution to obtain a mixture that is 7% alcohol?
Answer by stanbon(75887)

How many liters of an 11% alcohol solution should be mixed with 60 liters of a 5% alcohol solution to obtain a mixture that is 7% alcohol?
---------
Amt of 5% solution = 60 L ; amt of active ingredient = 0.05*60 = 3 L
-------
Amt of 11% solution = x L ; amt of active ingredient = 0.11x L
---------------
Amt of 7% mixture = (60+x) L ; Amt of active ingre = 0.07(60+x) = 4.2+0.07x L
===================
EQUATION:
active + active = active
3 + 0.11x = 4.2+0.07x
0.04x = 1.2
x = 30 liters (amt of 11% solution that must be mixed)
==========
Cheers,
Stan H.